OpenStudy (anonymous):

Will fan and medal! Solve right triangle ABC: C= 90*, A= 15*, and side b= 17.6. I found that angle B= 75* but how do I use tangent and cosine to find the sides of the triangle?

OpenStudy (anonymous):



OpenStudy (cwrw238):

sin B = opposite side/ hypotenuse sin 75 = 17.6 / h h * sin 75 = 17.6 h = 17.6 / sin 75 now grab ur calculator and work it out

OpenStudy (cwrw238):

Note I could have let hypotenuse = c you can find the length of the other side a by using tan 15 = a / 17.6

OpenStudy (anonymous):

i got 18.220860…. is that right?

OpenStudy (anonymous):

the was just an three dots.

OpenStudy (cwrw238):

18.22 is ok - correct to 2 decimal places

OpenStudy (anonymous):

I think i was supposed to use tangent..the answer wouldn't have came out the same if i would have used tangent, right?

OpenStudy (anonymous):

or was that for side c??

OpenStudy (cwrw238):

to find the hypotenuse you use sine or cosine

OpenStudy (cwrw238):

that was for side c

OpenStudy (anonymous):

oh okay! that makes sense.

OpenStudy (cwrw238):

c = hypotenuse

OpenStudy (anonymous):

gottcha

OpenStudy (cwrw238):

i used h by really it was better to use c (opposite angle C)

OpenStudy (cwrw238):

did you want to use the cosine to find c not the sine?

OpenStudy (anonymous):

no, that's fine. I just forgot i could use sine as well for hypotenuse. what about side a? How would I find that?

OpenStudy (cwrw238):

to find side a you can use the tangent tan A = opposite / adjacent tan 15 = a / 17.6 a = 17.6 * tan 15

OpenStudy (cwrw238):

do you know SOH-CAH-TOA - helps you remember what each trig ratio means

OpenStudy (anonymous):

Oh yeah, now i feel silly for not seeing that one! hahaha

OpenStudy (anonymous):

Yes! Thank you so much for your help!

OpenStudy (cwrw238):

yw